# CityLens: Benchmarking Large Language-Vision Models for Urban Socioeconomic Sensing

### v1

@@ 3.3 Evaluation of Direct and Normalized Estimation
3.3.1 Overall Performance
We evaluate the performance of large language-vision models on all 11 tasks using both the Direct Metric Prediction and Normalized Estimation settings. To ensure robustness of the analysis, any data point for which either the simple or normalized $R^{2}$ is less than or equal to 0.5 is excluded, as such values generally reflect weak predictive performance and are not analytically meaningful. The final comparison is visualized in Figure 4. A few tasks such as House Price, Public Transport, and Building Height achieve relatively better $R^{2}$ scores under certain models and settings, e.g., House Price consistently exceeds 0.2 under the Direct setting. These tasks are likely more visually grounded, with cues such as building density, road layout, and commercial signage that can be directly observed from urban imagery. This suggests that some socioeconomic indicators may be approximated more easily when the visual-structural link is strong. However, the majority of results fall into the low or even negative $R^{2}$ range, indicating that the model’s predictions often fail to explain the variance in the ground-truth indicator values. This suggests that,
@@ 3.3.1 Overall Performance
while large language-vision models have made rapid progress in perception and reasoning, predicting region-level socioeconomic indicators remains a highly challenging and underexplored task. The models may still lack the necessary numerical grounding, contextual interpretation, and semantic alignment required to associate urban visual content with structured socioeconomic quantities. Even with normalization, which alleviates the demand for precision by coarsening the prediction space, performance remains weak across most tasks. In many cases, the model predictions tend to collapse toward city-wide averages or exhibit a narrow output range, suggesting a lack of sensitivity to fine-grained regional variation. This behavior indicates that the models may struggle to differentiate subtle socio-spatial differences across urban regions, especially when visual cues are weak or ambiguous. These findings highlight the inherent difficulty of the CityLens benchmark; models struggle not only when asked to predict exact values, but also when the task is simplified into normalized estimation. We also incorporate the Feature-Based Regression method into a unified evaluation. A comprehensive comparison across all three paradigms is visualized in Appendix Figure 9.
3.3.2 Task Preference Between Direct and Normalized Estimation
In Figure 4, the diagonal line indicates equal performance under both methods; points above it suggest that the task benefits more from normalization, while points below indicate a preference for direct estimation. This result highlights that different tasks tend to favor different estimation strategies, depending on the nature of the indicator and its visual and semantic properties.
@@ 3.3.2 Task Preference Between Direct and Normalized Estimation
Specifically, tasks such as Violent Crime, GDP, and Population are more frequently observed above the diagonal, suggesting that these indicators with limited direct visual correspondence benefit from a normalized formulation that emphasizes relative ranking rather than precise value prediction. These tasks are difficult to estimate accurately, but models may still capture coarse ordinal relationships across regions, aided by their global knowledge priors and implicit ranking sense. Conversely, tasks like Bachelor Ratio, House Price, Public Transport, and Accessibility to Health tend to fall below the diagonal, indicating better performance under the direct estimation setting. These tasks are often associated with clearer, more stable visual correlates, such as building types, infrastructure visibility, and environmental layout, which can support more precise image-to-value mappings. In addition, some indicators, e.g., Life Expectancy, exhibit narrower value ranges or lower variance, making them more amenable to direct value prediction. Moreover, for tasks like House Price and Bachelor Ratio, LLVMs may leverage latent knowledge about typical value scales across different cities, enabling surprisingly accurate numerical predictions.
Taken together, these findings emphasize the importance of task-specific method selection in socioeconomic indicator prediction. Some indicators benefit from reducing output precision to focus on broader distinctions, while others retain enough visual signal to support direct prediction. The CityLens benchmark thus not only tests model capacity, but also reveals the nuanced interplay between task semantics and prediction strategy.
3.4 Detailed Analysis
@@
3.4.1 City-Level Performance Variations
To better understand the variation in socioeconomic prediction outcomes across different cities, we conduct a city-level analysis for the GDP task under the Feature-Based Regression paradigm. Each of the 14 cities is represented by 100 regions, with Gemma3-12B extracting 13 visual features per street view image. The GDP task is selected because it is a global-scale indicator, available across the largest number of cities in CityLens. Among the 13 cities evaluated in the GDP prediction task, we observe considerable variation in model performance in Figure 5(a). Cities such as Shanghai, San Francisco, and Sao Paulo achieve $R^{2}$ scores above 0.43, indicating relatively strong predictive performance. One possible explanation for the strong performance in cities like Shanghai lies in their well-structured urban design and high alignment between street-level appearance and economic development. These cities tend to have clear visual stratification between affluent and less affluent areas, consistent architectural patterns and homogeneous zoning that make features more learnable and high quality, diverse street view coverage. In contrast, cities like Mumbai and Moscow yield near-zero or even negative $R^{2}$, which may be attributed to two key factors. First, there may be a weak alignment between street-level visuals and actual economic activity, especially in cities with spatially mixed development, where wealth and poverty coexist within the same region, blurring the visual economic signal. Second, the quality and coverage of street view images can be a limiting factor. Inconsistent image sources, low resolution, or sparse sampling reduce the availability of reliable visual cues, hindering feature extraction and degrading downstream prediction.
3.4.2 Impact of Satellite Imagery
In this part, we evaluate the impact of satellite imagery by comparing model performance in two configurations: with and without satellite imagery. We test 3 tasks: House price, Public transport, and Drive ratio using Gemini-2.0-Flash under the Direct Metric Prediction setting. As shown in Figure 5(b), the results reveal minimal performance differences between the two configurations, which appears counterintuitive given the common assumption that satellite imagery can provide high-level spatial cues, such as regional development intensity, land use, and urban rural gradient, which are potentially informative for socioeconomic prediction. One possible explanation is that street view imagery already provides sufficiently rich and fine-grained visual information, like building conditions and public infrastructure. These ground-level features may be more directly interpretable by current large language-vision models and more tightly coupled with the target indicators in these tasks. As a result, the one additional satellite image may offer redundant or less impactful information, leading to negligible performance gains.
3.4.3 Effect of Street View Image Quantity
@@
To evaluate the impact of street view image quantity on prediction performance, we conduct experiments using Llama4-Maverick on the House Price task under the Direct Metric Prediction setting. In this experiment, each region is represented by one satellite image and a varying number of street view images: 1, 5, 10, 15, or 20. The results are shown in Figure 5(c). In addition, we test a no-image baseline following the design in manvi2024geollm , where only the geographic coordinates and address are provided in the prompt without any visual input. In this setting, most responses from the model consist of refusals, indicating that the model lacks sufficient information to generate a confident estimate. Interestingly, in several cases, the model proactively suggests alternatives, such as checking local housing websites, demonstrating the LLVM’s conversational safety and reasoning abilities, but also its limitation in open-world knowledge retrieval. From the results, we observe a clear trend: increasing the number of street view images consistently improves model performance. As more images are included, $R^{2}$ increases and nRMSE decreases. This suggests that a richer visual context helps the model form a more accurate understanding of the region’s socioeconomic condition.
4 Related Works
4.1 Urban Socioeconomic Sensing
@@
A growing number of studies have attempted to predict socioeconomic indicators in urban environments, using diverse technical strategies. lbkgzhou and liu2023knowcl employ knowledge graph based approaches to support socioeconomic inference. Li et al. li2022predicting propose a contrastive learning framework based on structural urban imagery to support socioeconomic indicator prediction. Fan et al. fan2023uvi extract urban visual features from street view imagery using a computer vision model, applying them to predict socioeconomic indicators at scale across U.S. cities. Recent approaches have extended this paradigm to multi-modal vision-language settings. UrbanCLIP yan2024urbanclip and UrbanVLP hao2025urbanvlp utilize contrastive learning on urban images while incorporating textual prompts generated by large language models. GeoLLM manvi2024geollm extracts latent geospatial knowledge from large language models through fine-tuning and prompt design. manvi2024large investigates geographical bias in LLMs via zero-shot socioeconomic prediction, while li2024can evaluates the performance of LLMs on socioeconomic tasks across region-level and city-level granularity, comparing three different paradigms. Different from these works, we introduce CityLens, the first benchmark to systematically evaluate the ability of large vision-language models to predict socioeconomic indicators using both street view and satellite imagery. CityLens spans a total of 17 cities from 6 continents, 11 indicators across 6 domains.
4.2 Benchmarking LLM and LLVM
@@
In recent years, LLMs have demonstrated strong commonsense and reasoning abilities, rapidly surpassing evaluation benchmarks across a range of domains that were previously established. Thus, developing diverse and challenging benchmarks to systematically assess the capabilities of LLMs has become an emerging and highly active research direction. A variety of benchmarks have been proposed, including conversational benchmarks such as Chatbot Arena chiang2024chatbot and MT-Bench bai2024mt , coding benchmarks such as SWE-Bench jimenez2023swe and LiveCodeBench jain2024livecodebench , mathematical benchmarks such as Gaokao-MathQA zhong2023agieval and SciBench wang2023scibench , as well as agent benchmarks like AgentBench liu2023agentbench and ToolBench qin2023toolllm . Furthermore, numerous multimodal benchmarks have also been introduced to evaluate the capabilities of Large Language-Vision Models, including comprehensive benchmarks like SEED-Bench li2023seed and MMMU yue2024mmmu , and domain-specific benchmarks such as CARES xia2024cares , UrBench zhou2025urbench , and DABench hu2024infiagent . While there are some benchmarks include urban imagery (e.g., CityBench feng2024citybench and UrBench zhou2025urbench ), they are not specifically designed for urban socioeconomic sensing tasks and address only a very limited subset of such tasks. Therefore, in this paper, after a comprehensive review of previous works in urban sensing, we propose a new benchmark to fill this gap and provide opportunities to bridge LLMs with urban sensing applications.
@@ 5 Conclusion
In this paper, we introduce CityLens, a benchmark for evaluating the ability of large language-vision models to predict socioeconomic indicators from satellite and street view imagery. Through extensive experiments across 3 evaluation paradigms and 17 state-of-the-art models, we find that while current models exhibit promising perceptual abilities on certain visually grounded tasks, they still face major challenges in making accurate and generalizable predictions across domains and regions. CityLens provides a foundation for analyzing these limitations and motivates further research into enhancing the capabilities of large language-vision models in urban socioeconomic sensing.
6 Discussion

### v2

@@ 3.2 Evaluation of Direct and Normalized Estimation
Overall Performance
We evaluate the performance of large vision-language models on all 11 tasks using both the Direct Metric Prediction and Normalized Estimation settings. To ensure meaningful analysis, we exclude model-task pairs with $R^{2}\leq-0.5$ under either setting and choose to abstract away the model identity. The final comparison is visualized in Figure 5, where each point represents the performance of a specific model on a specific task, evaluated under both estimation settings. A few tasks such as House Price, Public Transport, and Building Height achieve relatively better $R^{2}$ scores under certain models and settings, e.g., House Price consistently exceeds 0.2 under the Direct setting. These tasks are likely more visually grounded, with cues such as building density, road layout, and commercial signage that can be directly observed from urban imagery. This suggests that some socioeconomic indicators may be approximated more easily when the visual-structural link is strong. However, the majority of results fall into the low or even negative $R^{2}$ range, indicating that the model’s predictions often fail to explain the variance in the ground-truth indicator values. This suggests that, the models may still lack the necessary numerical grounding, contextual interpretation, and semantic alignment required to associate urban visual content with structured socioeconomic quantities. Even with normalization, which alleviates the demand for precision by coarsening the prediction space, performance remains weak across most tasks. In many cases, the model predictions tend to collapse toward city-wide averages or exhibit a narrow output range, suggesting a lack of sensitivity to fine-grained regional variation. This behavior indicates that the models may struggle to differentiate subtle socio-spatial differences across urban regions, especially when visual cues are weak or ambiguous.
@@
These findings highlight the inherent difficulty of CityLens; models struggle not only when asked to predict exact values, but also when the task is simplified into normalized estimation. We also incorporate the Feature-Based Regression method into a unified evaluation. A comprehensive comparison across all three paradigms is visualized in Appendix Figure 12.
Task Preference Between Direct and Normalized Estimation
@@
In Figure 5, the diagonal line indicates equal performance under both methods; points above it suggest that the task benefits more from normalization, while points below indicate a preference for direct estimation. This result highlights that different tasks tend to favor different estimation strategies, depending on the nature of the indicator and its visual and semantic properties. Specifically, tasks such as Violent Crime, GDP, and Population are more frequently observed above the diagonal, suggesting that these indicators with limited direct visual correspondence benefit from a normalized formulation that emphasizes relative ranking rather than precise value prediction. These tasks are difficult to estimate accurately, but models may still capture coarse ordinal relationships across regions, aided by their global knowledge priors and implicit ranking sense. Conversely, tasks like Bachelor Ratio, House Price, Public Transport, and Accessibility to Health tend to fall below the diagonal, indicating better performance under the direct estimation setting. These tasks are often associated with clearer, more stable visual correlates, such as building types, infrastructure visibility, and environmental layout, which can support more precise image-to-value mappings. In addition, some indicators, e.g., Life Expectancy, exhibit narrower value ranges or lower variance, making them more amenable to direct value prediction. Moreover, for tasks like House Price and Bachelor Ratio, LVLMs may leverage latent knowledge about typical value scales across different cities, enabling surprisingly accurate numerical predictions. Taken together, these findings emphasize the importance of task-specific method selection in socioeconomic indicator prediction. The CityLens benchmark thus not only tests model capacity, but also reveals the nuanced interplay between task semantics and prediction strategy.
3.3 Influence of Geographic Variation and Input Composition
City-Level Performance Variations
To better understand the variation in socioeconomic prediction outcomes across different cities, we conduct a city-level analysis for the GDP task under the Feature-Based Regression paradigm. Each of the 13 cities is represented by 100 regions, with Gemma3-12B extracting 13 visual features per street view image. Among the 13 cities evaluated in the GDP prediction task, we observe considerable variation in model performance in Figure 6(a). Cities such as Shanghai, San Francisco, and Sao Paulo achieve $R^{2}$ scores above 0.43, indicating relatively strong predictive performance. One possible explanation for the strong performance in cities like Shanghai lies in their well-structured urban design and high alignment between street-level appearance and economic development. These cities tend to have clear visual stratification between affluent and less affluent areas, consistent architectural patterns and homogeneous zoning that make features more learnable and high quality, diverse street view coverage. In contrast, cities like Mumbai and Moscow yield near-zero or even negative $R^{2}$, which may be attributed to two key factors. First, there may be a weak alignment between street-level visuals and actual economic activity, especially in cities with spatially mixed development, where wealth and poverty coexist within the same region, blurring the visual economic signal. Second, the quality and coverage of street view images can be a limiting factor. Inconsistent image sources, low resolution, or sparse sampling reduce the availability of reliable visual cues, hindering feature extraction and degrading downstream prediction.
Impact of Input Modalities
In this part, we evaluate the impact of input modalities by comparing model performance in three configurations: using both, only street view, and only satellite imagery. We test House price, Public transport, and Drive ratio using Gemini-2.0-Flash under the Direct Metric Prediction setting. Contrary to prior findings that satellite imagery is often more discriminative than street view imagery for urban representation (Sun et al., 2025; Hao et al., 2025), our results in Figure 6(b) show that using street view images alone achieves performance comparable to using both street view and satellite imagery, and significantly outperforms using satellite imagery alone. This suggests that street view images provide more semantically rich and fine-grained visual cues, such as building façades, commercial signage, and infrastructure quality. These ground-level features are likely more tightly coupled with socioeconomic indicators and more readily interpreted by current LVLMs, which have been pretrained extensively on image–language pairs featuring such localized, human-centric content. While satellite imagery exhibits weaker predictive performance, it still contributes independent spatial context, such as urban morphology and building layout. However, it may not offer the same level of semantic density at the resolution used in CityLens.
@@ 3.3 Influence of Geographic Variation and Input Composition
Effect of Street View Image Quantity
@@
To evaluate the impact of street view image quantity on prediction performance, we conduct experiments using Llama4-Maverick on the House Price task under the Direct Metric Prediction setting. Each region includes one satellite image and a varying number of street view images: 1, 5, 10, 15, or 20. We also test a no-image baseline following the design in Manvi et al. (2024b), where only the geographic coordinates and address are provided. In this setting, the model often refuses to respond, occasionally suggesting external resources like local housing websites, demonstrating both its conversational safety and limitation in open-world knowledge retrieval. From Figure 6(c), we observe a clear trend: increasing the number of street view images consistently improves model performance. This suggests that a richer visual context helps the model form a more accurate understanding of the region’s socioeconomic condition.
3.4 Reasoning Capabilities and Model Architecture Comparison
Tasks HP PT DR BR
Gemma3-12B -0.145 0.226 0.120 -0.019
Gemma3-12B-CoT 0.121 0.156 0.076 -0.049
Llama4-Maverick 0.795 0.459 0.406 0.503
Llama4-Maverick-CoT 0.794 0.392 0.270 0.167
Gemini-2.0-Flash 0.373 0.395 0.561 0.177
Gemini-2.0-Flash-CoT 0.602 0.436 0.508 0.222
((a))
((b))
((c))
Figure 7: (a) Model performance with and without CoT prompting. (b) Performance of reasoning models. (c) Comparison of different vision encoders used within the evaluated models.
Comparison of Chain-of-Thought Prompting vs. Standard Prompting
Following the designs of Zhang et al. (2025b) and Xu et al. (2024), we implement a Chain-of-Thought (CoT) prompting strategy tailored to the urban socioeconomic sensing context. The example CoT prompt is shown in A.7.4. We conduct an evaluation of CoT prompting on four representative tasks using three different models under the Direct Metric Prediction setting. As the Figure 7(a) below shows, we observe that the effect of CoT prompting varies by task. For the House Price task, CoT almost consistently improves performance across all models, suggesting it helps with the complex reasoning involved in interpreting housing-related visual and semantic cues. In contrast, for the Drive Ratio task, CoT often reduces performance, possibly because this task relies more on direct visual features rather than step-by-step reasoning. From a model perspective, Gemini-2.0-Flash benefits most consistently from CoT, with improvements across nearly all tasks. However, Llama4-Maverick shows performance drops with CoT. One possible explanation is that Llama4-Maverick already possesses strong internal reasoning abilities, and the externally imposed CoT structure may not align with its learned inference patterns, leading to performance degradation.
Reasoning Model Performance with Standard Prompt
We test three advanced reasoning models on three tasks that are inherently difficult to predict directly from images under the Normalized setting. We note that CoT prompting and reasoning models represent two distinct approaches: CoT focuses on prompt design, injecting explicit reasoning steps into the input to guide model thinking, while reasoning models are evaluated using standard prompts to assess their intrinsic reasoning capabilities. As shown in Figure 7(b), none of the powerful reasoning models achieve strong performance across all tasks, highlighting the challenge of predicting abstract social indicators from visual data alone. For instance, GLM-4.1v-Thinking performs best on Violent Crime ($R^{2}$=0.145) but poorly on Bachelor Ratio, while o4-mini achieves the highest result on Bachelor Ratio ($R^{2}$=0.465) but returns a negative $R^{2}$ on Violent Crime. Interestingly, we observe that Gemini-2.5-Flash performs better in its non-thinking version compared to its thinking version. This may be because urban socioeconomic sensing is not a purely logical reasoning task (as in math or code), but rather requires a nuanced combination of visual understanding and contextual inference. In such settings, reasoning-specific adaptations may not always align well with the nature of the task.
The Impact of Different Vision Encoders
We also follow the setup in Karamcheti et al. (2024) and conduct experiments using four models, which differ only in their vision encoders, under the Feature-Based Regression setting to investigate the role of visual backbones. This result in Figure 7(c) suggests that LVLMs initialized with CLIP as the vision encoder produce most informative and semantically aligned outputs for urban socioeconomic sensing. The improved downstream performance may stem from CLIP’s ability to extract visual cues that the language model can more effectively reason over. DINOv2 performs moderately but inconsistently, struggling on tasks like Life Expectancy, suggesting its self-supervised features may lack the semantic depth needed for abstract urban indicators. In contrast, SigLIP and IN1K perform consistently poorly, indicating that general-purpose contrastive or classification-based encoders are less effective at capturing relevant visual cues.
Figure 8: Representative error cases from Feature-Based Regression and CoT Prompting.
3.5 Error Analysis and Upper Bounds of LVLMs
Error Cases on Challenging Tasks
To better understand why LVLMs underperform on challenging tasks, we analyze representative errors observed under both the Feature-Based Regression and CoT Prompting setups. As illustrated in Figure 8, errors can arise from both visual perception and linguistic reasoning. For example, during feature extraction, Gemma3-4B fails to detect small but meaningful elements such as street signs, hallucinates non-existent persons, and underestimates visible greenery, assigning a low score to “Grass & Shrubs”. These errors reveal a lack of fine-grained visual grounding and semantic alignment, which can propagate into downstream reasoning and prediction. We also observe reasoning errors in the CoT setting. For instance, in a Mental Health prediction case, Gemma3-12B overly focuses on a few old and modest houses while overlooking numerous well-maintained, even upscale beachfront apartments visible in the street view images. Moreover, it fails to leverage the region’s proximity to water, which is a known factor with strong aesthetic and calming effects that are directly linked to mental well-being. This suggests that current LVLMs may struggle to appropriately weigh holistic environmental cues during reasoning.
The Potential of Fine-Tuned LVLMs
We conduct preliminary supervised fine-tuning on Qwen2.5-VL-7B, Qwen3-VL-8B, and Llama3.2-VL-11B, using additional CityLens data not included in the benchmark under the Direct Metric Prediction setup. Since the training data lacks samples for House Price, Violent Crime, and Life Expectancy, these tasks are excluded from evaluation. While general state-of-the-art LVLMs often perform poorly on CityLens benchmark, frequently yielding near-zero or even negative $R^{2}$ scores, our results in Table 2 show that fine-tuned LVLMs, regardless of base model or parameter size, achieve consistently strong performance across nearly all tasks. These findings highlight the promising potential of LVLMs for urban socioeconomic sensing, and further provide a preliminary estimate of the upper bound that such models can achieve when properly adapted for this domain. This reinforces the central motivation behind CityLens and underscores the value of developing domain-specific LVLMs for addressing this socially important challenge.
Table 2: Results of Fine-Tuned LVLMs on the CityLens Benchmark.
Tasks GDP Pop. PT DR BH MH AH BR
Fine-tuned Qwen2.5-VL-7B 0.628 0.231 0.502 0.628 0.872 0.418 0.364 0.442
Fine-tuned Qwen3-VL-8B 0.626 0.107 0.545 0.638 0.869 0.397 0.304 0.536
Fine-tuned Llama3.2-VL-11B 0.562 0.287 0.348 0.256 0.829 0.248 0.256 0.157
4 Related Work
Urban Socioeconomic Sensing
@@ 4 Related Work
A growing number of studies have attempted to predict socioeconomic indicators in urban environments. Zhou et al. (2023) and Liu et al. (2023b) employ knowledge graph based approaches to support socioeconomic inference.  Li et al. (2022) propose a contrastive learning framework based on structural urban imagery to support socioeconomic prediction.  Fan et al. (2023) extract features from street view imagery via a computer vision model to predict urban indicators. More recent studies have begun incorporating LLMs into this domain (Liu et al., 2025; Ouyang et al., 2024)Yan et al. (2024) and Hao et al. (2025) combine contrastive learning on urban images with LLM-generated textual prompts. Manvi et al. (2024b) extracts geospatial knowledge from LLMs through fine-tuning and prompt design, while Li et al. (2024b) evaluates LLMs on socioeconomic tasks across region-level and city-level granularity. Different from these works, CityLens is the first benchmark to systematically evaluate the ability of LVLMs to predict socioeconomic indicators using both street view and satellite imagery.
Benchmarking LLM and LVLM
@@
In recent years, LLMs have rapidly advanced in commonsense and reasoning, leading to the creation of diverse benchmarks across domains. These include benchmarks for code (Jimenez et al., 2023; Jain et al., 2024), mathematics (Zhong et al., 2023; Wang et al., 2023), city (Feng et al., 2025a), as well as agent-based tasks  (Liu et al., 2023a; Qin et al., 2023). Furthermore, numerous multimodal benchmarks have also emerged for LVLMs, including comprehensive benchmarks like Li et al. (2023) and Yue et al. (2024), and domain-specific ones such as Xia et al. (2024)Zhou et al. (2025), and Hu et al. (2024). While there are some benchmarks include urban imagery (e.g., Feng et al. (2025c; b)Zhang et al. (2025a)Zhou et al. (2025)), they are not specifically designed for urban socioeconomic sensing and address only a very limited subset of such tasks. Therefore, after a comprehensive review of previous works, we propose a new benchmark to fill this gap and provide opportunities to bridge LVLMs with urban sensing applications.
@@ 5 Conclusion
In this paper, we introduce CityLens, a benchmark for evaluating the ability of large vision-language models to predict socioeconomic indicators from satellite and street view imagery. Through extensive experiments across 3 evaluation paradigms and 17 state-of-the-art models, we find that while current models exhibit promising perceptual abilities on certain visually grounded tasks, they still face major challenges in making accurate and generalizable predictions across domains and regions. CityLens provides a foundation for analyzing these limitations and motivates further research into enhancing the capabilities of large vision-language models in urban socioeconomic sensing.
Acknowledgements
This work was supported in part by the National Key Research and Development Program of China under Grant 2024YFC3307602, in part by the Guangdong Provincial Talent Program under Grant No.2023JC10X009, in part by the National Key Research and Development Program of China under grant 2024YFC3307603. This work was also sponsored by Tsinghua-Toyota Joint Research Institute Inter-disciplinary Program.
Ethics statement
Privacy
All street view images in CityLens are sourced from platforms such as Google Street View, Baidu Maps, and Mapillary, which enforce automatic blurring of sensitive visual information, including faces and license plates. No personal or identifiable data is collected, stored, or annotated by the authors. All images are used exclusively for academic research purposes. Furthermore, the image resolution is coarse-grained, and all imagery is de-identified, ensuring that no individual-level visual content is exposed. The socioeconomic indicators used in the benchmark are aggregated at the regional level (e.g., Census Tract) rather than the individual level, further mitigating privacy risks.
Geographic Bias and Fairness
CityLens covers 17 cities across all six continents, ensuring a high level of geographic diversity. However, for certain indicators, particularly those culturally sensitive data, ground-truth labels are unavailable in some cities, most notably in regions within the Global South. As a result, the distribution of prediction tasks is uneven across regions, which may raise concerns of underrepresentation of Global South cities. Beyond task availability, such geographic imbalances in data coverage may also contribute to potential geographic biases in how current LVLMs generalize across diverse socioeconomic and cultural contexts. We include a preliminary bias audit in Appendix A.9, where we observe noticeable differences in performance across cities. We highlight this as both a diagnostic insight and an opportunity for future research, particularly in addressing fairness and robustness in cross-regional prediction scenarios.
Misuse Declaimer
CityLens is designed solely for research and evaluation purposes. It should not be used to inform real-world decisions in areas such as policing, health, or public resource allocation. The benchmark includes sensitive socioeconomic and crime-related indicators that are highly context-dependent. Any use of model outputs evaluated on CityLens for operational or policy purposes must be preceded by ethical review, fairness assessment, and domain-specific validation. We strongly discourage the use of this benchmark for surveillance or enforcement without appropriate safeguards.
Reproducibility statement
Our work aims to ensure transparency and reproducibility through full release of both data and code. All resources are made publicly available via https://github.com/tsinghua-fib-lab/CityLens and https://huggingface.co/datasets/Tianhui-Liu/CityLens-Data.
Dataset
The CityLens Benchmark includes two versions of the dataset:
  • CityLens (Google/Baidu-based): This version uses street view images retrieved via Google and Baidu APIs. Due to licensing restrictions, we cannot directly distribute the images. Instead, we follow the practice of Fan et al. (2023); Huang et al. (2024); Wang et al. (2025), and provide a complete list of pano ids along with download scripts for automated image retrieval. Additionally, we release the task data covering 11 socioeconomic indicators in the repository.
  • CityLens-Mapillary (Open-source): This version leverages street view images from Mapillary, an open-source street-level imagery platform. All images in this version are fully accessible.
Both versions of the dataset share the same set of satellite images.
Code
We release all code required for data processing and evaluation. Detailed instructions and usage examples are provided in the README.md file within the repository.
Appendix A Appendix
A.1 The Use of Large Language Models
This paper made limited use of large language models to assist with improving the clarity and stylistic quality of writing. The use of LLMs was restricted to language editing and polishing of author-written content. No text was generated solely by the model without human verification, and all technical claims, experimental results, and interpretations were entirely conceived, written, and validated by the authors. The authors retained full responsibility for the content of the paper and ensured that the use of LLMs complied with relevant ethical standards and publication guidelines.
A.2 Discussion